# Hypergraphs as Weighted Directed Self-Looped Graphs: Spectral Properties, Clustering, Cheeger Inequality

### v1

@@ 4.2 Approximate Linear Optimality of HyperClus-G
From Theorem 8, and the relaxation of the restriction, the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $L_{sym}$ is an approximation of both optimal NCut and the NCut of the returned cluster. Therefore, the returned cluster is approximately linearly optimal in terms of NCut. We then prove that it is also approximately linearly optimal in terms of conductance.
Using the $\leq$” side of Hypergraph Cheeger Inequality (Theorem 10), $\lambda$ satisfies $\lambda\leq 2\Phi(\mathcal{H})$, where $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$ is the optimal conductance. Additionally, we have the following lemma which connects NCut and conductance.
Lemma 11.
@@
From Definition 10 and 12,
$\begin{split}\Phi(\mathcal{S})&=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}\\ &=\frac{\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}\\ &\leq(\frac{1}{vol(\mathcal{S})}+\frac{1}{vol(\bar{\mathcal{S}})})\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}\\ &=NCut(\mathcal{S},\bar{\mathcal{S}})\end{split}$ (42)
@@
which shows the approximately linear optimality of the returned partition from HyperClus-G.
5 Algorithm Complexity
5.1 Worst-case Time Complexity of HyperClus-G
Step 5, computing the eigenvector associated with the second smallest eigenvalue takes $O(|\mathcal{V}|^{3})$. And step 6 takes $O(|\mathcal{V}|)$. Therefore step 5 and step 6 together take $O(|\mathcal{V}|^{3})$.
Therefore the worst-case complexity of HyperClus-G is
5.2 Worst-case Space Complexity of HyperClus-G
We make the same assumption as in Section 5.1. The storage of $R,W,D_{\mathcal{V}},D_{\mathcal{E}}$ takes $O(m)$. The storage of $P$ and $L$ takes $O(|\mathcal{V}|^{2})$. The storage of stationary distribution and the intermediate results takes $O(|\mathcal{V}|)$. The intermediate results for eigendecomposition take $O(|\mathcal{V}|^{2})$.
Therefore, the worst-case space complexity for HyperClus-G is
@@ 6 Supportive Experiments
In this section, we demonstrate the effectiveness of our HyperClus-G on real-world data. We first describe the experimental settings and then discuss the experiment results. Details to reproduce the results are provided in Appendix B.
@@ 6.1 Global Partitioning Setup
Datasets and Hypergraph Constructions. We use 9 datasets, all from UC Irvine Machine Learning Repository. The datasets cover biology, computer science, business, and other subject areas. The construction progress of hypergraphs is the same as the seminal works [52, 33, 24]: for each dataset, each instance is a vertex, and we use a group of hyperedges to capture each feature. For example, in the Mushroom dataset, for its “stalk-shape” feature, we connect all the instances that have an enlarging stalk shape by a hyperedge, and connect all the instances that have a tapering stalk shape by another hyperedge. We use one hyperedge for each categorical feature. For numerical features, we first quantize them into bins of equal size, then map them to hyperedges. More details of each dataset and construction can be found in Appendix B.2.
@@
In other words, for any hyperedge $e$ and any vertex $v$ in class $k$, $\gamma_{e}(v)$ is the number of vertices in class $k$ that $e$ connects. This assignment makes the vertex weights for different hyperedges highly different. We intentionally and implicitly encode the label information into the vertex weights to validate that our HyperClus-G can take the fine-grained information that is usually ignored by other hypergraph modeling methods.
Settings and Metrics. Among the 9 datasets, 5 of them have 2 classes, and 4 of them have at least 3 classes. For 2-class datasets, we apply HyperClus-G to partition the whole EDVW hypergraph into 2 clusters. For k-class datasets ($k\geq 3$), we call HyperClus-G iteratively for $k-1$ times to get a k-way clustering. We first measure the quality of clusters by NCut. The 2-way clustering NCut has been provided in Definition 10, and we further define the k-way clustering NCut here.
Definition 13.
For any k-way clustering of vertex set $\mathcal{V}$,
Another metric on cluster quality is whether the partition aligns with the ”ground-truth” label data. The vertices in the same cluster are determined by the algorithms to be closely related in structure, while the vertices in the same label class should also be internally similar. These two similarities usually align, and therefore we greedily match the clusters with classes and see if each matching has a high F1 score. For 2-way clustering, we report the F1 scores of both clusters. For k-way clustering, we report the weighted F1 scores of all clusters. More details are provided in Appendix B.3.
Baselines. We compare our HyperClus-G with multiple state-of-the-art methods. STAR++ and CLIQUE++ expansions convert the hypergraph into weighted graphs and then adopt spectral clustering on graphs. Moreover, since it is proven that, the EIVW Laplacian proposed by [52] is equal to the Laplacian of the corresponding STAR++ expanded graph [1, 7], our STAR++ baseline is equivalent to spectral clustering on EIVW hypergraphs by ignoring the assigned EDVW. DiffEq [41] is based on differential operators. node2vec [21] and event2vec (a.k.a., hyperedge2vec) [16] are two unsupervised graph embedding methods. We first convert the hypergraph into the required input graph form, then call the k-means algorithm after we get the vertex embedding. We also compare with the NCut of the actual labeled classes ($\mathcal{V}_{1},...,\mathcal{V}_{k}$ partitioning). For the nondeterministic baselines, we run the experiment 10 times and report the mean. The standard deviations are relatively small, and we put them in Appendix B.6.1. More details on baselines are provided in Appendix B.5.
@@ 6.2 Results
The global partitioning experiment results are shown in Tables 3 and 4. The best results are marked in bold, and the second-best results are marked with underlines. For 2-way clustering, we also consider that the two F1s of a high-quality partition should be fairly similar. First, our algorithm HyperClus-G generally outperforms the baseline methods in terms of NCut. It is worth noting that, facing a combinatorial optimization problem, our HyperClus-G constantly outperforms the actual-class partition, and may already get very close to the optimal NCut. Second, STAR++ and CLIQUE++ expansions, as graph approximations for hypergraphs, are very strong baselines and can generate sub-optimal clustering. Third, in terms of F1 scores matched with actual classes, our HyperClus-G also achieves the best performance in general. The only exception is Digit k-way clustering, where HyperClus-G achieves the second best, but is very close to the best. This again shows that STAR++ and CLIQUE++ expansions are strong baselines. Unsupervised embedding methods can beat STAR++ and CLIQUE++ expansions on smaller datasets. Fourth, though HyperClus-G is not designed for k-way clustering, applying it multiple times returns a high-quality k-way clustering. Fifth, it turns out that the second smallest eigenvalue of $L_{sym}$ is a good approximation for NCut (Refer to Section 6.3 for details), which is consistent with our theoretical analysis. Sixth, we report the execution time in Table 5. in fact, for NCut in Rice, Covertype, and F1 in Digit, where STAR++ or CLIQUE++ expansion achieves the best performance, it needs to take 10×10$\times$10 × time compared to HyperClus-G.
Table 5: Execution Time($\downarrow$) on Global Partitioning Task in seconds. Full Table in Appendix B.6.2.
@@
relative error $\frac{|eigenvalue-ncut|}{ncut}$ 3.397% 24.91% 7.993% 2.994% 2.783%
@@ 6.3 Validation of the Eigenvalue Approximation
According to our theoretical study of Theorem 8, and the later transformation of the problem, the second smallest eigenvalue of the symmetric normalized random-walk-based hypergraph Laplacian $L_{sym}$ should be an approximation of the NCut value by relaxing the combinatorial optimization constraint. We validate this by showing the 2-way NCut value of the clustering result from our algorithm HyperClus-G, and compare it with the second smallest eigenvalue of $L_{sym}$. The data are shown in Table 6. It turns out that the two numbers are positively correlated and the error tends to be small. The Rice dataset, which has much more hyperedges than other datasets, has the largest relative error.
@@ A.1 Definition of Hyperpath and Hypergraph Connectivity
[52] (Hyperpath and hypergraph connectivity). We say there is a $hyperpath$ between vertices $v_{1}$ and $v_{k}$ when there is a sequence of distinct vertices and hyperedges $v_{1},e_{1},v_{2},e_{2},...,e_{k-1},v_{k}$ such that $v_{i}\in e_{i}$ and $v_{i+1}\in e_{i}$ for $11\leq i\leq k-$ 1. A hypergraph $\mathcal{H}$ is connected if there exists a hyperpath for any pair of vertices.
@@ A.3 $P$ is Irreducible
First, we create a new matrix $Q$ by replacing $P_{u,v}$ by 1 if $P_{u,v}>0$, and prove that $Q$ is irreducible, hence $P$ holds the same feature.
By the assumption that the undirected hypergraph $H$ is connected, every vertex $u$ is reachable from any vertex $v\neq u$ through a sequence of hyperedges. $Q_{u,v}=1$ if and only if $P_{u,v}>0$.
By definition of $P$,
$P_{u,v}=\sum_{e\in E(v)}\frac{\omega(e)}{d(u)}\frac{\gamma_{e}(v)}{\delta(e)}>0$
if there is a hyperedge connecting $v$ and $u$. Therefore, the directed graph $G$ with $Q$ as the adjacency matrix can be constructed through replacing every hyperedge by a directed fully-connected clique.
For any vertex pair $v,u$, by the assumption of $H$, there exists a sequence of hyperedges from $v$ to $u$, hence in $G$ there also exists a sequence of directed edges from $v$ to $u$. This means $G$ is strongly connected and $Q$ is irreducible. According to Theorem 6.2.44 in Matrix Analysis [25], $P$ is irreducible. ∎
@@ A.5 Additional Lemma to Support Random-walk-based Hypergraph Laplacian (Definition 5)
Therefore for regular graphs, if we follow the same definition of hypergraph Laplacian as of equation 52,
@@
which has the same eigenvectors as the regular graph Laplacian $D-A$.
Appendix B Experiemnt Details
B.1 Environments
@@ B.2 Datasets
Mushroom (https://archive.ics.uci.edu/dataset/73/mushroom) includes categorical descriptions of 8124 mushrooms in the Agaricus and Lepiota Family. Each species is labeled as edible or poisonous. This dataset contains missing data in the ”stalk-root” feature. As [52] did, we removed the feature that contains missing labels. The two clusters have 4208 and 3916 instances, respectively.
Rice (https://archive.ics.uci.edu/dataset/545/rice+cammeo+and+osmancik), a.k.a., Rice (Cammeo and Osmancik). A total of 3810 rice grain’s images were taken for the two species, processed, and feature inferences were made. 7 morphological features were obtained for each grain of rice. This dataset does not have missing data. The feature types are real, including integer values, continuous values, and binary values. The two clusters have 2180 and 1630 instances, respectively. For each continuous feature, we first find the maximum value of this feature, then normalize all the numbers in this feature by the maximum value. This results in 10 bins of equal size $[0,0.1],(0.1,0.2],...(0.9,1]$. Then we convert the quantified bins to categorical features by using a categorical feature to mark which bins a continuous value originally belongs to.
Car (https://archive.ics.uci.edu/dataset/19/car+evaluation), a.k.a., Car Evaluation. Originally, this dataset contained 1728 cars with labeled acceptability related to 6 features. This dataset does not have missing data. We extract all the cars that are labeled ”good” or ”vgood” (very good) and construct a smaller dataset of 134 cars. The two clusters have 65 and 69 instances, respectively.
@@
Covertype (https://archive.ics.uci.edu/dataset/31/covertype). The task of this dataset is the classification of pixels into 7 forest cover types based on attributes such as elevation, aspect, slope, hillshade, soil-type, and more. We follow [24] and extract the instances of classes 4 and 5 to construct a dataset for 2-way clustering. The numerical feature values in this dataset can vary within a large range. Therefore, we first quantize the numerical values into 10 bins of equal size. We use the same strategy as described in the Rice dataset above to convert the features into categorical features. The two clusters have 9493 and 2747 instances, respectively.
@@
Hypergraph $|\mathcal{V}|$ $|\mathcal{E}|$ $\sum_{v\in\mathcal{V}}E(v)$ classes $|\mathcal{E}|$ clique # original features Subject Are Feature Type
@@
B.3 Global Partitioning Metrics

### v2

@@ 4.2 Approximate Linear Optimality of HyperClus-G
Optimality in terms of NCut. From Theorem 1, and the relaxation of the restriction, the second smallest eigenvector $\lambda$ of the normalized hypergraph Laplacian $L_{sym}$ is an approximation of both optimal NCut and the NCut of the returned cluster. Therefore, the returned cluster is approximately linearly optimal in terms of NCut.
Optimality in terms of conductance. Using the RHS of Hypergraph Cheeger Inequality (Theorem 3), $\lambda$ satisfies $\lambda\leq 2\Phi(\mathcal{H})$, where $\Phi(\mathcal{H})=\mathop{\min}_{\mathcal{S}\subseteq\mathcal{V}}{\Phi(\mathcal{S})}$ is the optimal conductance. Additionally, we have the following corollary from Definition 16 and 18, connecting NCut and conductance.
Corollary 20.
@@
From Definition 16 and 18,
$\begin{split}\Phi(\mathcal{S})&=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}=\frac{\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}\\ &\leq(\frac{1}{vol(\mathcal{S})}+\frac{1}{vol(\bar{\mathcal{S}})})\sum_{u\in\mathcal{S},v\in\bar{\mathcal{S}}}\phi(u)P_{u,v}=NCut(\mathcal{S},\bar{\mathcal{S}})\end{split}$ (26)
@@
which shows the approximately linear optimality of the returned partition from HyperClus-G. In fact, we can make the $\approx"$ more rigorous by observing that $\lambda=\mathop{\min}_{x\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}}\frac{1}{2}R(x)\leq\frac{1}{2}R(\tilde{x})=NCut(\mathcal{S}_{return},\bar{\mathcal{S}}_{return})$, where $\tilde{x}$ is defined according to Theorem 1 with $\mathcal{S}=\mathcal{S}_{return}$. However, it is impossible to bound $NCut(\mathcal{S}_{\mathrm{return}},\bar{\mathcal{S}}_{\mathrm{return}})$ in terms of big-$O$ of $\lambda$, since there exist counterexamples for sign-threshold spectral graph partitioning [^3] For example, dumbbell graphs consisting of two large cliques (or expanders) connected by a single edge. (and thus also for EDVW hypergraphs) where $\lambda$ can be arbitrarily small while $NCut$ remains arbitrarily large.
@@ 5 Supportive Experiments
In this section, we demonstrate the effectiveness of our HyperClus-G on real-world data. We first describe the experimental settings and then discuss the experiment results. Details are provided in Appendix C.
@@ 5.1 Global Partitioning Setup
Datasets and Hypergraph Constructions. We use 9 datasets, all from the UC Irvine Machine Learning Repository. The datasets cover biology, computer science, business, and other subject areas. The construction progress of hypergraphs is the same as the seminal works (Zhou et al., 2006; Li & Milenkovic, 2018; Hein et al., 2013): for each dataset, each instance is a vertex, and we use a group of hyperedges to capture each feature. For example, in the Mushroom dataset, for its “stalk-shape" feature, we connect all the instances that have an enlarging stalk shape by a hyperedge, and connect all the instances that have a tapering stalk shape by another hyperedge. We use one hyperedge for each categorical feature. For numerical features, we first quantize them into bins of equal size, then map them to hyperedges. More details of each dataset and construction can be found in Appendix C.2.
@@
In other words, for any hyperedge $e$ and any vertex $v$ in class $k$, $\gamma_{e}(v)$ is the number of vertices in class $k$ that $e$ connects. This assignment makes the vertex weights for different hyperedges highly different. We intentionally and implicitly encode the label information into the vertex weights to validate that our HyeprClus-G can take the fine-grained information that is usually ignored by other hypergraph modeling methods.
Settings and Metrics. Among the 9 datasets, 5 of them have 2 classes, and 4 of them have at least 3 classes. For 2-class datasets, we apply HyperClus-G to partition the whole EDVW hypergraph into 2 clusters. For k-class datasets ($k\geq 3$), we call HyperClus-G iteratively for k1k-1 times to get a $k$-way clustering. We first measure the quality of clusters by NCut. The 2-way clustering NCut has been provided in Definition 16, and we further define the $k$-way clustering NCut here.
@@
Definition 21.
For any $k$-way clustering of vertex set $\mathcal{V}$,
Another metric on cluster quality is whether the partition aligns with the "ground-truth" label data. The vertices in the same cluster are determined by the algorithms to be closely related in structure, while the vertices in the same label class should also be internally similar. These two similarities usually align, and therefore we greedily match the clusters with classes and see if each matching has a high F1 score. For 2-way clustering, we report the F1 scores of both clusters. For $k$-way clustering, we report the weighted F1 scores of all clusters. More details are provided in Appendix C.3.
Baselines. We compare our HyperClus-G with multiple state-of-the-art methods. STAR++ and CLIQUE++ expansions convert the hypergraph into weighted graphs and then adopt spectral clustering on graphs. Moreover, since it is proven that, the EIVW Laplacian proposed by (Zhou et al., 2006) is equal to the Laplacian of the corresponding STAR++ expanded graph (Agarwal et al., 2006; Chitra & Raphael, 2019), our STAR++ baseline is equivalent to spectral clustering on EIVW hypergraphs by ignoring the assigned EDVW. DiffEq (Takai et al., 2020) is based on differential operators. node2vec (Grover & Leskovec, 2016) and event2vec (a.k.a., hyperedge2vec) (Fu et al., 2019) are two unsupervised graph embedding methods. We first convert the hypergraph into the required input graph form, then call the k-means algorithm after we get the vertex embedding. We also compare with the NCut of the actual labeled classes ($\mathcal{V}_{1},...,\mathcal{V}_{k}$ partitioning). For the nondeterministic baselines, we run the experiment 10 times and report the mean. The standard deviations are relatively small, and we put them in Appendix C.6.1. More details on baselines are provided in Appendix C.5.
@@ 5.2 Results
The global partitioning experiment results are shown in Tables 4 and 5. The best results are marked in bold, and the second-best results are marked with underlines. For 2-way clustering, we also consider that the two F1s of a high-quality partition should be fairly similar. First, our algorithm HyperClus-G generally outperforms the baseline methods in terms of NCut. It is worth noting that, facing a combinatorial optimization problem, our HyperClus-G constantly outperforms the actual-class partition, and may already get very close to the optimal NCut. Second, STAR++ and CLIQUE++ expansions, as graph approximations for hypergraphs, are very strong baselines and can generate sub-optimal clustering. Third, in terms of F1 scores matched with actual classes, our HyperClus-G also achieves the best performance in general. The only exception is Digit $k$-way clustering, where HyperClus-G achieves the second best, but is very close to the best. This again shows that STAR++ and CLIQUE++ expansions are strong baselines. Unsupervised embedding methods can beat STAR++ and CLIQUE++ expansions on smaller datasets. Fourth, though HyperClus-G is not designed for $k$-way clustering, applying it multiple times returns a high-quality $k$-way clustering. Fifth, it turns out that the second smallest eigenvalue of $L_{sym}$ is a good approximation for NCut (Refer to Section 5.3 for details), which is consistent with our theoretical analysis. Sixth, we report the execution time in Table 6. in fact, for NCut in Rice, Covertype, and F1 in Digit, where STAR++ or CLIQUE++ expansion achieves the best performance, it needs to take $10\times$ time compared to HyperClus-G.
Table 6: Execution Time($\downarrow$) on Global Partitioning Task in seconds. Full Table in Appendix C.6.2.
@@
relative error |eigenvaluencut||eigenvalue-ncut| / ncut{ncut} 3.397% 24.91% 7.993% 2.994% 2.783%
@@ 5.3 Validation of the Eigenvalue Approximation
According to our theoretical study of Theorem 1, and the later transformation of the problem, the second smallest eigenvalue of the symmetric normalized random-walk-based hypergraph Laplacian $L_{sym}$ should be an approximation of the NCut value by relaxing the combinatorial optimization constraint. We validate this by showing the 2-way NCut value of the clustering result from our algorithm HyperClus-G, and compare it with the second smallest eigenvalue of $L_{sym}$. The data are shown in Table 7. It turns out that the two numbers are positively correlated, and the error tends to be small. The Rice dataset, which has much more hyperedges than other datasets, has the largest relative error.
@@ 6 Conclusion
Acknowledgment
This work is supported by National Science Foundation under Award No. IIS-2117902 and AFOSR (FA9550-24-1-0002). The views and conclusions are those of the authors and should not be interpreted as representing the official policies of the funding agencies or the government.
@@ A.1 Definition of Hyperpath and Hypergraph Connectivity
(Zhou et al., 2006) (Hyperpath and hypergraph connectivity). We say there is a hyperpathhyperpath between vertices $v_{1}$ and $v_{k}$ when there is a sequence of distinct vertices and hyperedges $v_{1},e_{1},v_{2},e_{2},...,e_{k-1},v_{k}$ such that $v_{i}\in e_{i}$ and $v_{i+1}\in e_{i}$ for $1\leq i\leq k-1$. A hypergraph $\mathcal{H}$ is connected if there exists a hyperpath for any pair of vertices.
@@ A.3 $P$ is Irreducible
First, we create a new matrix $Q$ by replacing $P_{u,v}$ by 1 if $P_{u,v}>0$, and prove that $Q$ is irreducible. Hence, $P$ holds the same feature.
By the assumption that the undirected hypergraph $H$ is connected, every node $u$ is reachable from any node $v\neq u$ through a sequence of hyperedges. $Q_{u,v}=1$ if and only if $P_{u,v}>0$.
By the definition of $P$,
$P_{u,v}=\sum_{e\in E(v)}\frac{\omega(e)}{d(u)}\frac{\gamma_{e}(v)}{\delta(e)}>0$ (32)
if there is a hyperedge connecting $v$ and $u$.
Therefore, the directed graph $G$ with $Q$ as the adjacency matrix can be constructed by replacing every hyperedge with a directed fully-connected clique.
For any node pair $v,u$, by the assumption of $H$, there exists a sequence of hyperedges from $v$ to $u$; hence, in $G$, there also exists a sequence of directed edges from $v$ to $u$. This means $G$ is strongly connected, and $Q$ is irreducible. According to Theorem 6.2.44 in Matrix Analysis (Horn & Johnson, 2012), $P$ is irreducible. ∎
@@ A.5 Additional Lemma to Support Random-walk-based Hypergraph Laplacian (Definition 9)
Therefore, for regular graphs, if we follow the same definition of hypergraph Laplacian as of equation 35,
@@
which has the same eigenvectors as the regular graph Laplacian DAD-A.
A.6 Proof of Theorem 14
Proof.
Notice that when $x^{T}Ax$ is a scalar,
$x^{T}Ax=(x^{T}Ax)^{T}=x^{T}A^{T}x$ (39)
substituting $A=\Pi P$, we have
$x^{T}\Pi Px=x^{T}P^{T}\Pi x$ (40)
Also note that
[…]
Then,
[…]
A.7 Proof of Theorem 1
Proof.
$x(u)=\begin{cases}\sqrt{\frac{vol(\bar{\mathcal{S}})}{vol(\mathcal{S})}}&\text{if }u\in\mathcal{S}\\ -\sqrt{\frac{vol(\mathcal{S})}{vol(\bar{\mathcal{S}})}}&\text{if }u\in\bar{\mathcal{S}}\end{cases}$ (43)
From the definition of NCutNCut in Eq 16, we can compute $R(x)$ as follows:
A.8 Proof of Theorem 19
Proof.
Recall from the definition 18, 12 and 13, the numerator and denominator are both non-negative, and
$\Phi(\mathcal{S})=\frac{|\partial\mathcal{S}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}=\frac{|\partial\bar{\mathcal{S}}|}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}$ (45)
Since $|\partial\mathcal{S}|\leq vol(\mathcal{S}),|\partial\bar{\mathcal{S}}|\leq vol(\bar{\mathcal{S}})$, $0\leq\Phi(\mathcal{S})\leq\frac{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}{\min(vol(\mathcal{S}),vol(\bar{\mathcal{S}}))}=1$. ∎
A.9 Proof of Theorem 3
Proof.
From Theorem 14,
Thus, the LFS has been proven. For the RHS, assume $w$ is the eigenvector of the normalized Laplacian associated with $\lambda$, let vector $f=\Pi^{-\frac{1}{2}}w$. Then,
Therefore, $\forall u$
[…]
One can validate that
Continue with Equation 50 with Theorem 3 in (Chung, 2005),
[…]
Using Equation 51, continue with Equation 53,
Appendix B Algorithm Complexity
B.1 Worst-case Time Complexity of HyperClus-G
Step 5, computing the eigenvector associated with the second smallest eigenvalue, takes $O(|\mathcal{V}|^{3})$. And step 6 takes $O(|\mathcal{V}|)$. Therefore step 5 and step 6 together take $O(|\mathcal{V}|^{3})$.
Therefore the worst-case complexity of HyperClus-G is
B.2 Worst-case Space Complexity of HyperClus-G
We make the same assumption as in Section B.1. The storage of $R,W,D_{\mathcal{V}},D_{\mathcal{E}}$ takes $O(m)$. The storage of $P$ and $L$ takes $O(|\mathcal{V}|^{2})$. The storage of stationary distribution and the intermediate results takes $O(|\mathcal{V}|)$. The intermediate results for eigendecomposition take $O(|\mathcal{V}|^{2})$.
Therefore, the worst-case space complexity for HyperClus-G is
Appendix C Experiemnt Details
C.1 Environments
@@ C.2 Datasets
Mushroom (https://archive.ics.uci.edu/dataset/73/mushroom) includes categorical descriptions of 8124 mushrooms in the Agaricus and Lepiota Family. Each species is labeled as edible or poisonous. This dataset contains missing data in the "stalk-root" feature. As (Zhou et al., 2006) did, we removed the feature that contains missing labels. The two clusters have 4208 and 3916 instances, respectively.
Rice (https://archive.ics.uci.edu/dataset/545/rice+cammeo+and+osmancik), a.k.a., Rice (Cammeo and Osmancik). A total of 3810 rice grain’s images were taken for the two species, processed, and feature inferences were made. 7 morphological features were obtained for each grain of rice. This dataset does not have missing data. The feature types are real, including integer values, continuous values, and binary values. The two clusters have 2180 and 1630 instances, respectively. For each continuous feature, we first find the maximum value of this feature, then normalize all the numbers in this feature by the maximum value. This results in 10 bins of equal size [0,0.1],(0.1,0.2],(0.9,1][0,0.1],(0.1,0.2],...(0.9,1]. Then, we convert the quantified bins to categorical features by using a categorical feature to mark which bins a continuous value originally belongs to.
Car (https://archive.ics.uci.edu/dataset/19/car+evaluation), a.k.a., Car Evaluation. Originally, this dataset contained 1728 cars with labeled acceptability related to 6 features. This dataset does not have missing data. We extract all the cars that are labeled "good" or "vgood" (very good) and construct a smaller dataset of 134 cars. The two clusters have 65 and 69 instances, respectively.
@@
Covertype (https://archive.ics.uci.edu/dataset/31/covertype). The task of this dataset is the classification of pixels into 7 forest cover types based on attributes such as elevation, aspect, slope, hillshade, soil-type, and more. We follow (Hein et al., 2013) and extract the instances of classes 4 and 5 to construct a dataset for 2-way clustering. The numerical feature values in this dataset can vary within a large range. Therefore, we first quantize the numerical values into 10 bins of equal size. We use the same strategy as described in the Rice dataset above to convert the features into categorical features. The two clusters have 9493 and 2747 instances, respectively.
@@
Hypergraph $|\mathcal{V}|$ $|\mathcal{E}|$ $\sum_{v\in\mathcal{V}}E(v)$ # classes $|\mathcal{E}|$ clique # original features Subject Are Feature Type
@@
C.3 Global Partitioning Metrics